Details.

Power up WFS from PDU.

Left power off the bayside stages.

Plataform installed.

Waiting to have the people finish and release plataform

* from HW gui.
  • 1x1 cropped.
  • Check ocam frames
  • OK.
  • cool down ocam2k.
* CCD47
  • 4x4 setup.
  • OK
* filter wheel.

* fov_select.py
  • Select #8

# name x y

1: Argos 64.403 21.503

2: ilocRetro 62.100 36.200

3: nomic 64.285 20.713

4: ilocArgos 59.505 24.001

5: nrm 60.072 26.926

6: ilocSky 62.101 35.800

7: imaging2 66.312 28.060

8: shark (*) 66.498 20.601

9: ales 63.667 19.543

10: nales 66.680 25.140

11: other 56.931 36.951

12: fizeau 65.407 17.802

13: retro 63.700 34.901

14: nulling 68.426 17.719

15: test 69.000 23.500

16: argos 66.002 21.508

17: agpm 62.366 24.631

18: imaging 67.433 23.789

Which one you want to set (enter number or name, -1 to exit):

8

File fov.txt has been updated with the "shark" configuration

fov_select_lbtiSX.png

JUan CArlos moved to another meeting. Starting at 9am

Jenny keep working on LBTI

Alessandro and Maria access to the platform. Stop AO part.

10:47 am waiting.

11:20 waiting.

12:15.PM
  • Shark-Nir done with the hardware. Released for use of Argos cal source.
  • Reconfiguring for SHARK-NIR.

  • Authorize Shark.
  • bayside stages: 60.5,20.6,62
Falt adsec: 541.

Rest/set

1-18 19:40:13.375304| MAIN > Maximum mode index used to apply flat -> # 641

OSS deploy.

Fri Nov 18 19:40:52.293 2022 OSS:sx.swa:deployArmAR: RPC: OK, CMD: success.

12:58 PM

Shark team at the telescope. Wait to have the clear.

In meantime setup Argos cal unit.
  • The OSS didn't work. Reset din't work
  • Argos operatred from the Argos interface.
  • Status before reset: Missing parameters.
  • Status after reset: SimotionOK

Start No disturbance. from start.
  • No PSF n the CCD47
  • defocus hexapod-> no PSF
  • Checking hexapod positons.

  • wiew vignetting i the CCD47. Right side of the pupl.
  • Adjustment of coma correction using the PSF.
This can be due to the deploy arm of the Shark-nir. The team has problems to control via soft.

1:14PM
  • Luca goes up to check the arm.
  • Corrected some aberrations
  • Stop at this point to allow theam to go to the telescope.
  • Filter 2 to OD3
  • chamber lights on
  • Alignment almost center.

CCd47_lbti_20221118_arm_deploy_vign.png

1:16PM

Hold WFS side.

Adsec initial setup for mag 9.2
  • KL_v20
  • REC_LBTI_IIR_B1_500M.FITS
  • LBTI_IIR_B1_1000
1:44PM back luca from telescope

PSF went offline/online

Argos source lmap interface showed from open shuttter to close. error state: overcurrent

1:59PM

back in operations.

The shark-nir arm vignetting from the top of the CCD47

The previous vignetting is form a position of tha bayside stage x 6 mm off.

positions x->66.5mm

2:10PM
  • PRESETAO
  • filter 1 timeoout: 21:06:51.29
  • after the tv dark.
  • repeated the dark tv and it was fine.
  • PresetAO
  • OK
  • FLAO GUI.
  • center pupil.
  • ok
  • close loop
  • OK.

No disturbance:
  • close loop
  • manual gain.
  • 3.95,0.65,0.4
  • A lot vibration.
  • TN: 20221118_213054
  • GOPT RUN.
  • With those gain the Shark-nir has a good PSF to work wth it.
2:30PM

Paul asked for time to reset the TCS.

Shark-Nir keeps working until 3:15PM after hand over to Paul.

Shark asked to move PSF moving the stage in X axe.

gain 0.1

move stage x

step 0.2

66.5

66.0

3:00PM
  • Juan Carlos to the meeting.
  • Jenny keeping working wth shark-nir team
new position fov changed.

3:19PM
  • handover to Paul.
  • OCAm set save.
  • done.
  • w_stop_all
  • PDU off steps:
  • 5
  • 8
  • 11 not problem.
  • 15
  • 16

SAve...

-- JuanCarlosGuerra - 18 Nov 2022
Topic revision: r3 - 18 Nov 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback